# Impact of the latest 22Ne+ reaction rates on nucleosynthesis in massive stars and galactic chemical evolution

### v1

@@ 4.3 Observational data
The simulated GCE was compared with observational data collected from STELLAB: a code with a database of spectra observations from individual stars within the Milky Way (see CĆ“tĆ© etĀ al., 2017). Most of STELLABās data are from the galactic disc, halo, and globular clusters within the Milky Way. In this work, we use the available stellar observations from the Milky Way disk to compare with our GCE simulations.
@@ 5 GCE Results
In Figure 12, we show the pre-SN GCE curves for Cu, Zn, Ga, and Ge calculated from the massive star yields using different 22Ne+ reaction rates, in comparison with observational data. Abundance evolution for each element is shown as [X/Fe] ratios555We remind that [X/Y] = (mass fractions are used for and ). The solar values are from Asplund etĀ al. (2009). Only the gas component (i.e., isotopes not trapped within stars) is considered.
The GCE curves shown in Fig. 12 are produced using the yields from the nucleosynthesis calculated with LO12 and OT20 rates including their uncertainties. Additional curves are calculated with other 22Ne+ rates, in particular the recommended rates from Adsley etĀ al. (2021) (Adsley-TAMU, hereafter AD21), from Wiescher etĀ al. (2023) (Notre-Dame (ND), hereafter ND23), and the original Ritter etĀ al. (2018b) configuration (hereafter CR18). Although not shown in the figure, we confirmed that the upper and lower limits of the GCE curves using the yields from ND23 and AD21 rates lie within the curves bounded by the LO12 and OT20 rates.
@@
A wealth of observational data is available for Cu and Zn; therefore, these elements have been studied in detail in previous GCE research (e.g., Matteucci etĀ al., 1993; Timmes etĀ al., 1995; Chiappini etĀ al., 1997; Mishenia etĀ al., 2002; Bisterzo etĀ al., 2005; Kobayashi etĀ al., 2006; Mishenina etĀ al., 2017; Kobayashi etĀ al., 2020; Lach etĀ al., 2020). The availability of observational data for Ga and Ge is instead scarce, with only a handful of metal-poor stars measured in the range considered for this work ([Fe/H]).
@@
The [X/Fe] values relative to the GCE calculation performed with yields using other rates typically evolve between the OT20 and LO12 limits. Except for Zn, all simulations are consistent with the solar values within the observational uncertainties. The inconsistency found in Zn is not surprising because of the strong primary Zn component that is not accounted for in these simulations (Pignatari etĀ al., 2010). As mentioned earlier, the source of the primary Zn production is still a matter of debate (Bisterzo etĀ al., 2005), with possible large contributions from hypernovae in the early Galaxy (Kobayashi etĀ al., 2020, and references therein). Thus, the Zn discrepancy shown in Figure 12 is not related to the current sensitivity study of the 22Ne reaction rates.
@@
The variation of [X/Fe] obtained using the OT20 and LO12 recommended rates is summarized in Table 2. Here we report the variation for all elements between Cu and Sr. In general, 0.2-0.5 dex variation was found for [X/Fe]. However, the impact on the GCE of As and other heavier elements should be re-evaluated in the future since the massive starās -process contribution for As and other heavier elements is expected to be smaller compared to AGB stars and/or explosive nucleosynthesis (Kappeler etĀ al., 2011).
@@ 5.2 Resonance Sensitivity
In this section, we investigate the difference in the LO12 and OT20 rates in more detail. We study the impact on the GCE of the differences in the two rates, focusing on the contributions from individual resonances. The of relevant resonances to the weak -process identified as having large uncertainties by Ota etĀ al. (2021) and by the recent review article Wiescher etĀ al. (2023) are summarized in Table 3. It should be noted that the adopted strengths and upper limits (denoted by āā) have substantially different meanings in stellar reaction rates as described below. The upper limits are not calculated using the specified .
While the adopted strength was used in the stellar reaction calculations, the upper limit was used to determine the median and probability distribution assuming the Porter-Thomas distribution (Longland etĀ al., 2010, 2012) in the present work. As such, the actual (e.g., median) strength used in stellar rate calculations is much smaller than the upper limit used as it is in e.g., Angulo etĀ al. (1999); Kappeler etĀ al. (1994); Massimi etĀ al. (2017). For example, for the =11.17 MeV resonance in LO12 has a median of 10-8 eV ( 4.3010-7 eV upper limit). Thus, the resonanceās impact on the stellar yield is negligibly small using either the LO12 or OT20 (5.4210-11 eV) rate.
@@
In Case 2, the for the =11.17 MeV resonance was increased to 660 neV, which is obtained by Talwar etĀ al. (2016) from their indirect measurement. Note that the existence of the resonance was invalidated by the latest direct measurement experiment by Shahina etĀ al. (2022), and an upper limit of 150 neV was determined. Thus, the median strength should be much smaller (10 neV). Still, we use 660 neV to probe the maximum impact of this resonance.
In Case 3, we increased the for the =11.11 MeV resonance to 13.5 neV (approximately the value adopted by Massimi etĀ al. (2017); see Ota etĀ al. (2021)). It should be noted that the resonance is dominated by (,n) channel and the is constrained to the same level as the =11.17 MeV resonance (60 neV) by direct measurements (Jaeger etĀ al., 2001). The upper limit used by LO12 (58 neV) reflects this value. We studied this resonance only for its potential contributions, as theorized by Massimi etĀ al. (2017). However, since this is the lowest possible resonance for () channel, the weak -process could be activated at lower temperature (0.3 GK) if the resonance strength is high.
@@
In Case 6, we decreased the OT20ās (,) reaction rate at 1 GK by a factor of 2. Since a number of resonances (primarily 11.8 MeV and above) contribute to the 22Ne(,)25Mg reaction, the relevance of the 22Ne(,)26Mg becomes negligible. In this temperature range, we estimated the summed impact of all the high energy resonances by simply applying a correction factor to the reaction rate.
@@
Given the difference in at =11.32 MeV is only a factor of 3.3, it is clear that this is the most impactful on s-process nucleosynthesis in massive stars among the remaining uncertain in the 22Ne + reaction. The recent direct measurement by Shahina etĀ al. (2024) reported the to be 100 (22) eV, which varies from OT20 by a factor of 2.4. Its impact on the GCE is therefore estimated to be 0.15 dex (Cu and Zn) and 0.3 dex (Ga and Ge), which would result in elemental abundances distributed between OT20 and LO12. Upcoming verification of their results by future measurements in underground facilities can considerably reduce the uncertainties in the GCE (Masha & LUNA Collaboration., 2022; Best etĀ al., 2025). If is unambiguously constrained with a precision of better than 10ā20% (or e.g., =10-20 eV; which are typical or slightly better experimental uncertainties reported for this resonance), it will reduce the uncertainties from this resonance on the GCE simulations to 0.05 dex for Cu and Zn, and 0.1 dex for Ga and Ge, respectively.
@@
For the of the =11.17 MeV resonance (Case 2), the GCE varies by 0.2ā0.4 dex when increasing the from 0 to 660 neV. Assuming a simple linearity, we could safely say that the GCE uncertainties will be 0.05 dex for all the elements considered if the upper limit was reduced 10 times from 660 neV. If future experiments set an upper limit of 60 neV on this resonance (the current limit: 150 neV by Shahina etĀ al. (2022)), rather small GCE uncertainties (0.05 dex) would be achievable. If the for the =11.11 MeV resonance (Case 3) was similarly constrained by reducing the upper limit (13.5 neV) 10 times, the GCE uncertainties would be 0.05 dex for all the elements considered. Therefore, a 1 neV upper limit would be enough for this resonance. Again, no experimental evidence has been reported that the =11.11 MeV resonance has a large -cluster strength. In addition, the direct measurements show no signature of a strong resonance near the corresponding energy. Nevertheless, it will be beneficial to constrain the upper limit of this to 1 neV in future direct measurements due to its potential impact and experimental feasibility. At underground facilities such measurements can be possible (Rapagnani etĀ al., 2022). It should also be mentioned that these upper limits may slightly change as the neutron productions compete among different resonances using limited amounts of 22Ne. For instance, if the dominant =11.32 MeV resonance has a higher value, the impact of the lower energy resonances could be reduced. Thus, even higher upper limits on the resonances may be acceptable.
@@ 5.3 Post-SN GCE
The purpose of the present work is to study the impact of uncertainties of the 22Ne+ rates on the s-process in massive stars; we have explored this topic in great detail in the previous sections. As we have seen in Section 4.1, the -process component found in some CCSN models by Ritter etĀ al. (2018b) is not affected by the 22Ne+ rates, and depends instead on the CCSN explosion setup. It is possible to roughly estimate the primary CCSN production of Cu to 6-16%, based on observations of metal-poor stars (Pignatari etĀ al., 2010). A similar contribution could be estimated for Ge (around 20%) and Ga, but the limited number of spectroscopic observations for these elements limits the ability to derive relevant constraints.
Using the same GCE framework adopted for the previous simulations, we have calculated a GCE model using the CCSN yield table provided by Ritter etĀ al. (2018b). Figure 14 shows the GCE curves of Cu, Zn, Ga, and Ge. As a reference, the pre-SN GCE scatter between the LO12 upper limit and OT20 lower limit (see section 5.1) is shown as green bands, with the pre-SN GCE using the CR18 rate. Within the range covered in the figure, the GCE model using CCSN yields for Ga and Ge does not appear to diverge significantly from the GCE models adopting the pre-SN yields, with the few observations available consistent with the results.
Simulations of Cu and Zn show instead some tension with observations (e.g., Mishenina etĀ al., 2017). The [Cu/Fe] trend is consistent with the solar abundances, but it is quite flat for different . Observations show instead solar ratios between ā and [Fe/H] -0.5, but then they start to decrease for lower . Using the same phenomenological considerations in Pignatari etĀ al. (2010), this would imply that in the GCE model adopting Ritter etĀ al. (2018b) CCSN yields, about 30-50% of the solar Cu would be made by primary CCSN components, assuming that 1/3 or 1/2 of the solar Fe is provided by massive stars. Therefore, in Ritter etĀ al. (2018b) standard CCSN yields the primary Cu production is overestimated by a factor between 2 and 8 with respect to observations, and this is directly noticeable in the CCSN abundance profiles from 12 and 15 stars at 0.006 (see e.g., Figures 1 and 2 for the 15 Mā model at Z=0.0001). An overestimation compared to observations is also obtained for the [Zn/Fe] ratios. This is interesting, since Zn has been historically under-produced by GCE simulations, leading to scenarios where a boosted Zn contribution from hypernovae has been claimed at low (Kobayashi etĀ al., 2020, and references therein). Our GCE simulations have no hypernovae, but we still see the huge potential impact of the -process on Zn, from the most internal 0.1-0.2 Mā in the CCSN ejecta.
@@ 6 Summary and future work
While 22Ne(,)25Mg and 22Ne(,)26Mg are among the most important reactions for s-process nucleosynthesis, their reaction rates still have significant uncertainties. Jayatissa etĀ al. (2020); Ota etĀ al. (2020) and Ota etĀ al. (2021) used indirect measurements to report weaker strengths of relevant resonances, such as the =11.32 MeV for the (,n) channel and 11.17 MeV for the (,) channel. The most recent direct measurement by Shahina etĀ al. (2024) obtained a more moderate value of for the =11.32 MeV resonance which is more consistent with other experiments within their experimental uncertainties (22%). Their measurements of the for the =11.17 MeV are consistent with our measurements, implying that this resonance probably does not contribute significantly to -process nucleosynthesis. As described in Wiescher etĀ al. (2023), there are remaining uncertainties in for other resonances below and above the energies of these two resonances.
@@
The uncertainties from each of these resonances affect abundance yields of theoretical stellar models and the GCE of the s-process elements. This work provides the first detailed and quantitative estimation of the impact of different 22Ne+ rates on the s-process in massive stars and on the GCE of its main nucleosynthesis products. In particular, we focus on the production of the element group Cu, Zn, Ga, and Ge with Z = 29ā32. For stellar nucleosynthesis simulations, we used the post-processing nucleosynthesis code, MPPNP, developed by the NuGrid Collaboration. For GCE simulations, we used the GCE code, OMEGA+. For the analysis, we considered 14 different 22Ne+ reaction rate combinations. For each reaction rate configuration, the nucleosynthesis in 20 massive star models (four initial stellar masses, 12, 15, 20, and 25 Mā, and five metalicities, Z=0.02, 0.01, 0.006, 0.001, and 0.0001) were calculated, for a total of 280 nucleosynthesis models.
@@
In comparison with observations, the simulated GCE reproduce their solar abundances except for Zn. Our simulations underproduce Cu at low partly because of the uncertainties of the GCE parameters used in the present work. The stellar models that are used can also impact the obtained final yields. For instance, Pumo etĀ al. (2010) discusses that convective overshooting during the He core burning influences how much 22Ne is available and the duration of neutron-capture conditions. Their work demonstrates that using different values for the overshooting parameter (10-5ā0.035; c.f., 0.022 used (in the present models Ritter etĀ al., 2018b)), the weak abundances could decrease/increase by a factor of 2ā3 in 15ā25 Mā stars. Furthermore, fast rotating massive stars may potentially increase the process yields compared to the non-rotating stars used in the present work, especially at low (Pignatari etĀ al., 2008; Frischknecht etĀ al., 2016; Roberti etĀ al., 2024a). As well as () cross sections, uncertainties of the other nuclear cross sections relevant to neutron economy in the process stellar environment such as neutron poison (e.g., 16O()17O) could also help explain the inconsistency with the observations. The ratio of 17O()/17O() cross sections influences the available neutrons for the weak process. As shown in Best etĀ al. (2013), the ratioās current uncertainties could vary the abundances of weak elements by an order of magnitude. Therefore, many uncertainties in the physics input remain that hinder comparisons between the GCE simulations and observational data. Nonetheless, we consider the relative differences in GCE by different 22Ne reaction rates reported in the present work are still valid.
@@
We obtain that using our recent 22Ne(,)25Mg and 22Ne(,)26Mg rates, the elements Cu, Zn, Ga, and Ge result in up to a factor of 1.5-3 (0.2-0.45 dex) reductions in their galactic mass fractions compared to GCE models using the rates by Longland etĀ al. (2012). Such reductions are almost exclusively explained by the decrease of at the =11.32 MeV resonance. It confirms that this resonance is undoubtedly the current dominant source of the 22Ne+ rateās uncertainties for the s-process in massive stars.
@@
As such, it is critical to unambiguously constrain the of the =11.32 MeV resonance with a precision better than 10ā20% in future experiments. Thanks to the worldwide ongoing/planned experimental efforts at underground nuclear physics facilities such as LUNA, this should be feasible within the next few years. Lowering the upper limits on the of the =11.17 MeV and of the =11.11 MeV to negligible levels (60 neV and 1 neV, respectively) will also be possible (Rapagnani etĀ al., 2022). These experiments will have 10-100 times lower background environment compared to the work by Jaeger etĀ al. (2001). We have demonstrated that well-constrained 22Ne reaction rates will result in much smaller uncertainties than the observational data. This will provide a crucial milestone to better constrain stellar evolution models of massive stars and the chemical enrichment history of our Galaxy.
Finally, we discussed the GCE of Cu, Zn, Ga, and Ge using CCSN yields. The primary explosive components observed in metal-poor stars are found in our nucleosynthesis simulations. Such nucleosynthesis products are located in the deepest 0.1-0.2 Mā ejected, and they depend on the CCSN explosion setup. We show that they are not affected by the 22Ne reaction rates, or by their current uncertainties.
We emphasize the importance of including complete CCSN yields to compare stellar and GCE simulations with observational results more consistently, as they complement the pre-SN nucleosynthesis production. In future work, we indeed expect to use the s-process elements discussed here to constrain CCSN models by comparisons of GCE and stellar results with galactic archaeology observations, applying similar calibration methods adopted in the past using Ni (Limongi & Chieffi, 2018) or Zn (Nomoto etĀ al., 2013) as reference elements. In turn, we will apply the newly-constrained CCSN models to study GCE for all the elements.
prin_2022rjlwhn. We acknowledges support from the ChETEC-INFRA project funded by the European Unionās Horizon 2020 Research and Innovation programme (Grant Agreement No 101008324). We thanks access to viper, the University of Hull HPC Facility. This research has used the Astrohub online virtual research environment (https://astrohub.uvic.ca), developed and operated by the Computational Stellar Astrophysics group (http://csa.phys.uvic.ca) at the University of Victoria and hosted on the Computed Canada Arbutus Cloud at the University of Victoria. SO was supported in part by the IReNA visiting fellowship. This work benefited from

### v2

@@ 4.3 Observational data
The simulated GCE was compared with observational data collected from STELLAB: a code with a database of spectra observations from individual stars within the Milky Way (see 2017APJ_Cote). Most of STELLAB’s data are from the galactic disc, halo, and globular clusters within the Milky Way. In this work, we use the available stellar observations from the Milky Way disk to compare with our GCE simulations.
To supplement the STELLAB data for Ga and Ge, we imported stellar abundance data from JINAbase (2018ApJS..238...36A) and added observations from Ivans2003, and from cowan:05, Sneden1998, and Placco2015 for Ga and Ge, respectively.
@@ 5 GCE Results
In Figure 12, we show the pre-SN GCE curves for Cu, Zn, Ga, and Ge calculated from the massive star yields using different 22Ne+ reaction rates, in comparison with observational data. Abundance evolution for each element is shown as [X/Fe] ratios555We remind that [X/Y] = (mass fractions are used for and ). The solar values are from Asplund2009. Only the gas component (i.e., isotopes not trapped within stars) is considered.
The GCE curves shown in Fig. 12 are produced using the yields from the nucleosynthesis calculated with LO12 and OT20 rates including their uncertainties. Additional curves are calculated with other 22Ne+ rates, in particular the recommended rates from Adsley2021 (Adsley-TAMU, hereafter AD21), from Wiescher2023 (Notre-Dame (ND), hereafter ND23), and the original Ritter2018 configuration (hereafter CR18). Although not shown in the figure, we confirmed that the upper and lower limits of the GCE curves using the yields from ND23 and AD21 rates lie within the curves bounded by the LO12 and OT20 rates.
@@
A wealth of observational data is available for Cu and Zn; therefore, these elements have been studied in detail in previous GCE research (e.g., Matteucci1993; Timmes1995; 1997ApJ...477..765C; Mishenia2002; Bisterzo2005; Kobayashi2006; mishenina:17; kobayashi:20; lach:20). The availability of observational data for Ga and Ge is instead scarce, with only a handful of metal-poor stars measured in the range considered for this work ([Fe/H]).
@@
The [X/Fe] values relative to the GCE calculation performed with yields using other rates typically evolve between the OT20 and LO12 limits. Except for Zn, all simulations are consistent with the solar values within the observational uncertainties. The inconsistency found in Zn is not surprising because of the strong primary Zn component that is not accounted for in these simulations (2010ApJ...710.1557P). As mentioned earlier, the source of the primary Zn production is still a matter of debate (Bisterzo2005), with possible large contributions from hypernovae in the early Galaxy (kobayashi:20, and references therein). Thus, the Zn discrepancy shown in Figure 12 is not related to the current sensitivity study of the 22Ne reaction rates.
@@
The variation of [X/Fe] obtained using the OT20 and LO12 recommended rates is summarized in Table 2. Here we report the variation for all elements between Cu and Sr. In general, 0.2-0.5 dex variation was found for [X/Fe]. However, the impact on the GCE of As and other heavier elements should be re-evaluated in the future since the massive star’s -process contribution for As and other heavier elements is expected to be smaller compared to AGB stars and/or explosive nucleosynthesis (Kappeler2011).
@@ 5.2 Resonance Sensitivity
In this section, we investigate the difference in the LO12 and OT20 rates in more detail. We study the impact on the GCE of the differences in the two rates, focusing on the contributions from individual resonances. The of relevant resonances to the weak -process identified as having large uncertainties by Ota2021 and by the recent review article Wiescher2023 are summarized in Table 3. It should be noted that the adopted strengths and upper limits (denoted by ””) have substantially different meanings in stellar reaction rates as described below. The upper limits are not calculated using the specified .
While the adopted strength was used in the stellar reaction calculations, the upper limit was used to determine the median and probability distribution assuming the Porter-Thomas distribution (Longland2010; Longland2012) in the present work. As such, the actual (e.g., median) strength used in stellar rate calculations is much smaller than the upper limit used as it is in e.g., Angulo1999; Kappeler1994; Massimi2017. For example, for the =11.17 MeV resonance in LO12 has a median of 10-8 eV ( 4.3010-7 eV upper limit). Thus, the resonance’s impact on the stellar yield is negligibly small using either the LO12 or OT20 (5.4210-11 eV) rate.
@@
In Case 2, the for the =11.17 MeV resonance was increased to 660 neV, which is obtained by Talwar2016 from their indirect measurement. Note that the existence of the resonance was invalidated by the latest direct measurement experiment by Shahina2022, and an upper limit of 150 neV was determined. Thus, the median strength should be much smaller (10 neV). Still, we use 660 neV to probe the maximum impact of this resonance.
In Case 3, we increased the for the =11.11 MeV resonance to 13.5 neV (approximately the value adopted by Massimi2017; see Ota2021). It should be noted that the resonance is dominated by (,n) channel and the is constrained to the same level as the =11.17 MeV resonance (60 neV) by direct measurements (Jaeger2001). The upper limit used by LO12 (58 neV) reflects this value. We studied this resonance only for its potential contributions, as theorized by Massimi2017. However, since this is the lowest possible resonance for () channel, the weak -process could be activated at lower temperature (0.3 GK) if the resonance strength is high.
@@
In Case 6, we decreased the OT20’s (,) reaction rate at 1 GK by a factor of 2. Since a number of resonances (primarily 11.8 MeV and above) contribute to the 22Ne(,)25Mg reaction, the relevance of the 22Ne(,)26Mg becomes negligible. In this temperature range, we estimated the summed impact of all the high energy resonances by simply applying a correction factor to the reaction rate.
@@
Given the difference in at =11.32 MeV is only a factor of 3.3, it is clear that this is the most impactful on s-process nucleosynthesis in massive stars among the remaining uncertain in the 22Ne + reaction. The recent direct measurement by Shahina2024 reported the to be 100 (22) eV, which varies from OT20 by a factor of 2.4. Its impact on the GCE is therefore estimated to be 0.15 dex (Cu and Zn) and 0.3 dex (Ga and Ge), which would result in elemental abundances distributed between OT20 and LO12. Upcoming verification of their results by future measurements in underground facilities can considerably reduce the uncertainties in the GCE (Masha2022; Best2025). If is unambiguously constrained with a precision of better than 10–20% (or e.g., =10-20 eV; which are typical or slightly better experimental uncertainties reported for this resonance), it will reduce the uncertainties from this resonance on the GCE simulations to 0.05 dex for Cu and Zn, and 0.1 dex for Ga and Ge, respectively.
@@
For the of the =11.17 MeV resonance (Case 2), the GCE varies by 0.2–0.4 dex when increasing the from 0 to 660 neV. Assuming a simple linearity, we could safely say that the GCE uncertainties will be 0.05 dex for all the elements considered if the upper limit was reduced 10 times from 660 neV. If future experiments set an upper limit of 60 neV on this resonance (the current limit: 150 neV by Shahina2022), rather small GCE uncertainties (0.05 dex) would be achievable. If the for the =11.11 MeV resonance (Case 3) was similarly constrained by reducing the upper limit (13.5 neV) 10 times, the GCE uncertainties would be 0.05 dex for all the elements considered. Therefore, a 1 neV upper limit would be enough for this resonance. Again, no experimental evidence has been reported that the =11.11 MeV resonance has a large -cluster strength. In addition, the direct measurements show no signature of a strong resonance near the corresponding energy. Nevertheless, it will be beneficial to constrain the upper limit of this to 1 neV in future direct measurements due to its potential impact and experimental feasibility. At underground facilities such measurements can be possible (Rapagnani2022). It should also be mentioned that these upper limits may slightly change as the neutron productions compete among different resonances using limited amounts of 22Ne. For instance, if the dominant =11.32 MeV resonance has a higher value, the impact of the lower energy resonances could be reduced. Thus, even higher upper limits on the resonances may be acceptable.
@@ 5.3 Post-SN GCE
The purpose of the present work is to study the impact of uncertainties of the 22Ne+ rates on the s-process in massive stars; we have explored this topic in great detail in the previous sections. As we have seen in Section 4.1, the -process component found in some CCSN models by Ritter2018 is not affected by the 22Ne+ rates, and depends instead on the CCSN explosion setup. It is possible to roughly estimate the primary CCSN production of Cu to 6-16%, based on observations of metal-poor stars (2010ApJ...710.1557P). A similar contribution could be estimated for Ge (around 20%) and Ga, but the limited number of spectroscopic observations for these elements limits the ability to derive relevant constraints.
Using the same GCE framework adopted for the previous simulations, we have calculated a GCE model using the CCSN yield table provided by Ritter2018. Figure 14 shows the GCE curves of Cu, Zn, Ga, and Ge. As a reference, the pre-SN GCE scatter between the LO12 upper limit and OT20 lower limit (see section 5.1) is shown as green bands, with the pre-SN GCE using the CR18 rate. Within the range covered in the figure, the GCE model using CCSN yields for Ga and Ge does not appear to diverge significantly from the GCE models adopting the pre-SN yields, with the few observations available consistent with the results.
Simulations of Cu and Zn show instead some tension with observations (e.g., mishenina:17). The [Cu/Fe] trend is consistent with the solar abundances, but it is quite flat for different . Observations show instead solar ratios between ⊙ and [Fe/H] -0.5, but then they start to decrease for lower . Using the same phenomenological considerations in 2010ApJ...710.1557P, this would imply that in the GCE model adopting Ritter2018 CCSN yields, about 30-50% of the solar Cu would be made by primary CCSN components, assuming that 1/3 or 1/2 of the solar Fe is provided by massive stars. Therefore, in Ritter2018 standard CCSN yields the primary Cu production is overestimated by a factor between 2 and 8 with respect to observations, and this is directly noticeable in the CCSN abundance profiles from 12 and 15 stars at 0.006 (see e.g., Figures 1 and 2 for the 15 M⊙ model at Z=0.0001). An overestimation compared to observations is also obtained for the [Zn/Fe] ratios. This is interesting, since Zn has been historically under-produced by GCE simulations, leading to scenarios where a boosted Zn contribution from hypernovae has been claimed at low (kobayashi:20, and references therein). Our GCE simulations have no hypernovae, but we still see the huge potential impact of the -process on Zn, from the most internal 0.1-0.2 M⊙ in the CCSN ejecta.
@@ 6 Summary and future work
While 22Ne(,)25Mg and 22Ne(,)26Mg are among the most important reactions for s-process nucleosynthesis, their reaction rates still have significant uncertainties. Jayatissa2020; OTA2020135256 and Ota2021 used indirect measurements to report weaker strengths of relevant resonances, such as the =11.32 MeV for the (,n) channel and 11.17 MeV for the (,) channel. The most recent direct measurement by Shahina2024 obtained a more moderate value of for the =11.32 MeV resonance which is more consistent with other experiments within their experimental uncertainties (22%). Their measurements of the for the =11.17 MeV are consistent with our measurements, implying that this resonance probably does not contribute significantly to -process nucleosynthesis. As described in Wiescher2023, there are remaining uncertainties in for other resonances below and above the energies of these two resonances.
@@
The uncertainties from each of these resonances affect abundance yields of theoretical stellar models and the GCE of the s-process elements. This work provides the first detailed and quantitative estimation of the impact of different 22Ne+ rates on the s-process in massive stars and on the GCE of its main nucleosynthesis products. In particular, we focus on the production of the element group Cu, Zn, Ga, and Ge with Z = 29–32. For stellar nucleosynthesis simulations, we used the post-processing nucleosynthesis code, MPPNP, developed by the NuGrid Collaboration. For GCE simulations, we used the GCE code, OMEGA+. For the analysis, we considered 14 different 22Ne+ reaction rate combinations. For each reaction rate configuration, the nucleosynthesis in 20 massive star models (four initial stellar masses, 12, 15, 20, and 25 M⊙, and five metalicities, Z=0.02, 0.01, 0.006, 0.001, and 0.0001) were calculated, for a total of 280 nucleosynthesis models.
@@
In comparison with observations, the simulated GCE reproduce their solar abundances except for Zn. Our simulations underproduce Cu at low partly because of the uncertainties of the GCE parameters used in the present work. The stellar models that are used can also impact the obtained final yields. For instance, pumo:10 discusses that convective overshooting during the He core burning influences how much 22Ne is available and the duration of neutron-capture conditions. Their work demonstrates that using different values for the overshooting parameter (10-5–0.035; c.f., 0.022 used (in the present models Ritter2018)), the weak abundances could decrease/increase by a factor of 2–3 in 15–25 M⊙ stars. Furthermore, fast rotating massive stars may potentially increase the process yields compared to the non-rotating stars used in the present work, especially at low (pignatari:08; frischknecht:16; roberti:24). As well as () cross sections, uncertainties of the other nuclear cross sections relevant to neutron economy in the process stellar environment such as neutron poison (e.g., 16O()17O) could also help explain the inconsistency with the observations. The ratio of 17O()/17O() cross sections influences the available neutrons for the weak process. As shown in best:13, the ratio’s current uncertainties could vary the abundances of weak elements by an order of magnitude. Therefore, many uncertainties in the physics input remain that hinder comparisons between the GCE simulations and observational data. Nonetheless, we consider the relative differences in GCE by different 22Ne reaction rates reported in the present work are still valid.
@@
We obtain that using our recent 22Ne(,)25Mg and 22Ne(,)26Mg rates, the elements Cu, Zn, Ga, and Ge result in up to a factor of 1.5-3 (0.2-0.45 dex) reductions in their galactic mass fractions compared to GCE models using the rates by Longland2012. Such reductions are almost exclusively explained by the decrease of at the =11.32 MeV resonance. It confirms that this resonance is undoubtedly the current dominant source of the 22Ne+ rate’s uncertainties for the s-process in massive stars.
@@
As such, it is critical to unambiguously constrain the of the =11.32 MeV resonance with a precision better than 10–20% in future experiments. Thanks to the worldwide ongoing/planned experimental efforts at underground nuclear physics facilities such as LUNA, this should be feasible within the next few years. Lowering the upper limits on the of the =11.17 MeV and of the =11.11 MeV to negligible levels (60 neV and 1 neV, respectively) will also be possible (Rapagnani2022). These experiments will have 10-100 times lower background environment compared to the work by Jaeger2001. We have demonstrated that well-constrained 22Ne reaction rates will result in much smaller uncertainties than the observational data. This will provide a crucial milestone to better constrain stellar evolution models of massive stars and the chemical enrichment history of our Galaxy.
Finally, we discussed the GCE of Cu, Zn, Ga, and Ge using CCSN yields. The primary explosive components observed in metal-poor stars are found in our nucleosynthesis simulations. Such nucleosynthesis products are located in the deepest 0.1-0.2 M⊙ ejected, and they depend on the CCSN explosion setup. We show that they are not affected by the 22Ne reaction rates, or by their current uncertainties.
We emphasize the importance of including complete CCSN yields to compare stellar and GCE simulations with observational results more consistently, as they complement the pre-SN nucleosynthesis production. In future work, we indeed expect to use the s-process elements discussed here to constrain CCSN models by comparisons of GCE and stellar results with galactic archaeology observations, applying similar calibration methods adopted in the past using Ni (limongi:18) or Zn (Nomoto2013) as reference elements. In turn, we will apply the newly-constrained CCSN models to study GCE for all the elements.
prin_2022rjlwhn. We acknowledges support from the ChETEC-INFRA project funded by the European Union’s Horizon 2020 Research and Innovation programme (Grant Agreement No 101008324). We thanks access to viper, the University of Hull HPC Facility. This research has used the Astrohub online virtual research environment (https://astrohub.uvic.ca), developed and operated by the Computational Stellar Astrophysics group (http://csa.phys.uvic.ca) at the University of Victoria and hosted on the Computed Canada Arbutus Cloud at the University of Victoria. SO was supported in part by the IReNA visiting fellowship. This work benefited from